12. 6x +5x +7x

prime
(3x +1)(2x +1)
(2x +7)(3x +1)
x(6x +5x + 7)

13. 4x +14x -30

2(2x -3)(x +5)
2(2x -5)(x +3)
(4x -6)((x +5)
(2x -3)(2x +10)

14. 4x -8/(6x -12x)

(2x -2)/(3x -3x)
2/3
4/(6x)
2/(3x)

15. (3x -9)/(4x +20) . (x +3x -10)/(6x -12x)

3/4
(x -5)/(4(x +5))
(x -3)/(8x)
-3/8

16. (x -4x -5)/(x +2x +1) (x -5)/(x +4)

(x -5)/((x +1)(x +4))
5
(x +1)(x +4)/(x +1)
(x +4)/(x +1)

17. Find the slope of a line passing through (-4, 6) and (-2, 5)

-1/2
-2
1/2
1/6

18. Find the equation of the line passing through (2, 4) and (-3, 8)

4x +5y = 28
4x -5y = -12
5x +4y = 26
5x -4y = -6

19.
x +y = 2
3x -2y = 6



(1, 1)
(2,0)
Inconsistent
Dependent

20.
2x y = 8
x + 2y = 14



(6, 4)
(2, -4)
(0, 7)
(8, 3)

21.
x -2y = 12
x +3y = 2



(10, -1)
(0, -6)
(1, 1/3)
(8, -2)

22. 16x4y12

2xy
2xy
4xy6
4xy4

23. 2/12

1/6
6 /6
12 /6
3 /3

24.
The hypotenuse of a right triangle is 10. One side is 2 less than the other. Find the length of each leg.



3, 5
5, 7
8, 6
4, 6

25. x +7x +10 = 0

5, 2
-5, 2
-5, -2
-2, 5

Answer by mangopeeler07(462)

12. 6x +5x +7x

prime
(3x +1)(2x +1)
(2x +7)(3x +1)
x(6x +5x + 7)

Answer: x(6x +5x + 7)
*take out the common term, x
------------------------------------------------------------------------------
13. 4x +14x -30

2(2x -3)(x +5)
2(2x -5)(x +3)
(4x -6)((x +5)
(2x -3)(2x +10)

Answer: 2(2x-3)(x+5)
*Take out a two first, then factor
-------------------------------------------------------------------------------
14. 4x -8/(6x -12x)

(2x -2)/(3x -3x)
2/3
4/(6x)
2/(3x)

Answer: 2/3x
*Factor everything using common terms/factors, then cancel, then reduce
--------------------------------------------------------------------------------
15. (3x -9)/(4x +20) . (x +3x -10)/(6x -12x)

3/4
(x -5)/(4(x +5))
(x -3)/(8x)
-3/8

Answer: x-3/8x
*Factor using common terms/factors, then cancel, then multiply, then reduce
--------------------------------------------------------------------------------
16. (x -4x -5)/(x +2x +1) (x -5)/(x +4)

(x -5)/((x +1)(x +4))
5
(x +1)(x +4)/(x +1)
(x +4)/(x +1)
Answer: x+4/x+1
*Same concept as number 15
--------------------------------------------------------------------------------
17. Find the slope of a line passing through (-4, 6) and (-2, 5)

-1/2
-2
1/2
1/6
Answer: -1/2
*Difference of the y's over the difference of the x's
--------------------------------------------------------------------------------
18. Find the equation of the line passing through (2, 4) and (-3, 8)

4x +5y = 28
4x -5y = -12
5x +4y = 26
5x -4y = -6

Answer: 4x+5y=28
*Difference of the y's over the difference of the x's to obtain slope
*plug that and two coordinates in y=mx+b to solve for b
*manipulate the equation so that x and y are both on one side
--------------------------------------------------------------------------------
19.
x +y = 2
3x -2y = 6

(1, 1)
(2,0)
Inconsistent
Dependent

Answer: (2,0)
*Plug and check
--------------------------------------------------------------------------------
20.
2x y = 8
x + 2y = 14

(6, 4)
(2, -4)
(0, 7)
(8, 3)

Answer:(6,4)
*Same concept as 19
--------------------------------------------------------------------------------
21.
x -2y = 12
x +3y = 2

(10, -1)
(0, -6)
(1, 1/3)
(8, -2)

Answer:(8,-2)
*Same concept as 19 and 20
--------------------------------------------------------------------------------
24.
The hypotenuse of a right triangle is 10. One side is 2 less than the other. Find the length of each leg.

3, 5
5, 7
8, 6
4, 6

Answer: 8,6
*a^2+b^2=c^2
*a+b more than c
--------------------------------------------------------------------------------
25. x +7x +10 = 0

5, 2
-5, 2
-5, -2
-2, 5

Answer: -5, -2
*Factor
*Set each expression equal to zero and solve